A motorcycle has a constant acceleration of 4.33 m/s2. Both the velocity and acceleration of the motorcycle point in the same direction. How much time is required for the motorcycle to change its speed from (a)44.7 to 54.7 m/s, and (b)74.7 to 84.7 m/s?


1
Expert's answer
2022-02-06T14:33:40-0500

Given:

"a=4.33\:\rm m/s^2"


The time required for the motorcycle to change its speed

"\Delta t=\frac{\Delta v}{a}"

(a)

"\Delta t=\frac{54.7-44.7}{4.33}=2.31\:\rm s"

(b)

"\Delta t=\frac{84.7-74.7}{4.33}=2.31\:\rm s"